Stanford issues bonds dated January 1, 2019, with a par value of $246,000. The bonds annual contract rate is 8%, and interesRequired 1 Required 2 Required 3 How much total bond interest expense will be recognized over the life of these bonds? TotalRequired 1 Required 2 Required 3 Prepare an effective interest amortization table for these bonds. (Round all amounts to the
Answer #1
1
Discount 12490 =246000-233510
2
Total interest expense over life of bonds
6 payments of $ 9840 59040
Par value at maturity 246000
Total repaid 305040
Less: Amount borrowed 233510
Total bond interest expense 71530
3
Semiannual Interest period end Cash interest paid Bond interest expense Discount amortization Unamortized Discount Carrying value
01/01/2019 12490 233510
06/30/2019 9840 11676 1836 10654 235346
12/31/2019 9840 11767 1927 8727 237273
06/30/2020 9840 11864 2024 6703 239297
12/31/2020 9840 11965 2125 4578 241422
06/30/2021 9840 12071 2231 2347 243653
12/31/2021 9840 12187 2347 0 246000
Total 59040 71530 12490
Workings:
Cash interest paid 9840 =246000*8%*6/12
Bond interest expense = Carrying value X 10% X 6/12
Bond interest expense:
06/30/2019 11676 =233510*10%*6/12
12/31/2019 11767 =235346*10%*6/12
06/30/2020 11864 =237273*10%*6/12
12/31/2020 11965 =239297*10%*6/12
06/30/2021 12071 =241422*10%*6/12
12/31/2021 12187 =243653*10%*6/12